Suppose you are the money manager of a $4.42 million investment fund. The fund consists of four stocks with the following investments and betas: Stock Investment Beta A $ 260,000 1.50 B 400,000 (0.50 ) C 1,460,000 1.25 D 2,300,000 0.75 If the market's required rate of return is 8% and the risk-free rate is 6%, what is the fund's required rate of return? Do not round intermediate calculations. Round your answer to two decimal places. %
Answer #1
Step 1:
We need to calculate weighted average beta of fund
Stock Value Wight Beta Weighted average beta
a b c=b/4420000 d e=c*d
A 260000 0.06 1.50 0.09
B 400000 0.09 -0.50 -0.05
C 1460000 0.33 1.25 0.41
D 2300000 0.52 0.75 0.39
Total 4420000 1.00 0.8462
Step 2:
Return on fund:
Rf 6.00%
Rm 8.00%
Beta 0.8462
R = Rf+ B(Rm-Rf)
Where,
Rf = Risk Free Return
B= Beta
Rm-Rf= Risk Premium
=0.06+0.8462*(0.08-0.06)
=0.06+0.8462*(0.02)
=7.69 %